OpenStudy (sleepyjess):

Use basic identities to simplify the expression. sin\(\theta\) cos\(\theta\) sec\(\theta\) csc\(\theta\) \(sec^2\theta\) \(csc^2\theta\) 1 \(tan^2\theta\)

OpenStudy (anonymous):

did you type it properly? O.o how can you simplify '1' if its already in simpliest form? or is it one of the choices and you have to simplify \(sin \theta\ cos \theta\ sec \theta\ csc \theta\)?

OpenStudy (anonymous):

\(\sf sinθ\ cosθ\ secθ\ cscθ\\sinθ\ cosθ\ (\frac{1}{cos \theta}) (\frac{1}{sin \theta})\) cancel \(\sf sinθ\) and \(\sf cos θ\), you'll left with one